OpenStudy (anonymous):

Find all intercepts and asymptotes: x^2-4x And can someone tell me how to find them in general?

OpenStudy (anonymous):

your intercetps will be when y equals zero. i assume the function is y=x^2-4x?

OpenStudy (anonymous):

therefore factor out an x and get x(x-4)=0 one x =0 and another is x-4=0 --> x=4

OpenStudy (anonymous):

so your intercepts are x=0 and x=4 your asymptotes will be there if there are places in the function is undefined. therefore with this particular problem there are not any asymptotes.